### v1

# AuditLLM: A Tool for Auditing Large Language Models Using Multiprobe Approach

Abstract.
As Large Language Models (LLMs) gain wider adoption in various contexts, it becomes crucial to ensure they are reasonably safe, consistent, and reliable for an application at hand. This may require probing or auditing them. Probing LLMs with varied iterations of a single question could reveal potential inconsistencies in their knowledge or functionality. However, a tool for performing such audits with simple workflow and low technical threshold is lacking. In this demo, we introduce ”AuditLLM,” a novel tool designed to evaluate the performance of various LLMs in a methodical way. AuditLLM’s core functionality lies in its ability to test a given LLM by auditing it using multiple probes generated from a single question, thereby identifying any inconsistencies in the model’s understanding or operation. A reasonably robust, reliable, and consistent LLM should output semantically similar responses for a question asked differently or by different people. Based on this assumption, AuditLLM produces easily interpretable results regarding the LLM’s consistencies from a single question that the user enters. A certain level of inconsistency has been shown to be an indicator of potential bias, hallucinations, and other issues. One could then use the output of AuditLLM to further investigate issues with the aforementioned LLM. To facilitate demonstration and practical uses, AuditLLM offers two key modes: (1) Live mode which allows instant auditing of LLMs by analyzing responses to real-time queries; (2) Batch mode which facilitates comprehensive LLM auditing by processing multiple queries at once for in-depth analysis. This tool is beneficial for both researchers and general users, as it enhances our understanding of LLMs’ capabilities in generating responses, using a standardized auditing platform.
Large Language Model, Auditing LLMs Tools
* Work does not relate to position at Amazon.
copyright: nonecopyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Large Language Model  Auditing LLMs Tools
1. Introduction
In the rapidly evolving landscape of artificial intelligence, Large Language Models (LLMs) have emerged as cornerstones of contemporary AI applications, driving innovations in natural language processing (NLP), automated content creation, and decision support systems (Shi et al., 2023). Their ability to understand, generate, and interact using human-like language has paved the way for unprecedented advancements across various domains. For example, LLMs have become increasingly integrated into critical sectors, such as healthcare (Gu et al., 2021; Thirunavukarasu et al., 2023; Yang et al., 2022), education (Dai et al., 2023; Meyer et al., 2023; Kasneci et al., 2023), legal (Liu et al., 2023) and financial services (Zhang et al., 2023; Huang et al., 2023). Given the mission critical nature of these applications, it is imperative to ensure the LLMs are reliable and robust. Recognizing this, the academic and tech communities have underscored the importance of rigorous auditing mechanisms to probe LLM performance. Previous research has laid a solid foundation for understanding the potential of LLM auditing, focusing on identifying biases (Suguri Motoki et al., 2023; Talat et al., 2022; Thakur, 2023), hallucinations (Chen et al., 2023; Sadat et al., 2023; Yang et al., 2023), inconsistency (Tam et al., 2023; Ye et al., 2023), reliability (Shen et al., 2023; Zhong and Wang, 2023) and failures in model responses (Rastogi et al., 2023).
Building on the existing foundational work  (Rastogi et al., 2023), we introduce a novel auditing tool, AuditLLM, which provides a general-purpose solution for auditing LLMs. An simple but effective approach for probing LLMs involves deploying multiple versions of the same query to identify any inconsistencies in the model’s output. To implement this audit technique effectively, a mechanism for crafting these questions or probes is essential. While humans can come up with different ways to ask the same question for this probing purpose, it is not scalable. Therefore, we adopt the technique described in  Shah (2024) to streamline probe generation using a different LLM.
AuditLLM offers two primary modes for probing: live mode (used for demo and quick testing purposes) and batch mode (used for practical auditing needs).
2. Related Work
The utility of LLMs in enhancing information retrieval necessitates a critical assessment of their risks (Bender et al., 2021), a concern widely acknowledged by the academic community (Blodgett et al., 2020). Auditing LLMs, which focuses on evaluating specific attributes to address these risks, has become a key area of research. Studies, such as  (Mökander et al., 2023), have proposed a structured frameworks for LLM auditing, targeting ethical and social concerns through evaluation of model design, pre-release testing, and application usage. Similarly, (Wang et al., 2023) have compared GPT-4 and GPT-3.5, analyzing toxicity, bias, and robustness, while tools like Adatest (Ribeiro and Lundberg, 2022) and AdaTest++ (Rastogi et al., 2023) utilizes human feedback to identify LLM failures in model responses.
These efforts reflect advancements in LLM auditing for responsible use, however, none of these approaches have specifically targeted LLM inconsistencies. Our proposed framework, AuditLLM, has been designed specifically to address this gap in auditing LLMs. Given that inconsistencies could be harbingers for issues of bias and hallucination, focusing on auditing an LLM for its inconsistencies allows researchers and developers to be broader in their auditing efforts.
3. AuditLLM Tool Demonstration
3.1. System Overview
AuditLLM  [^1] Implemented using the Gradio Library in a HuggingFace space and available as a free, open-source project. serves as a systematic tool to reveal inconsistencies and conduct audits through the utilization of two LLMs: LLM1 and LLM2. As depicted in Figure 1, LLM1 is tasked with the creation of probes based on the given the prompt template, whereas LLM2 is responsible for generating responses to these probes. The prompt template for LLM1 is iteratively created using human-in-the-loop technique described by Shah (2024).
Figure 1. The AuditLLM system employs two LLMs: LLM1 generates probes and LLM2 provides responses to these probes.
AuditLLM introduces two distinct modes, live and batch, both designed to use multiple probes generated by LLM1 to detect discrepancies within the responses produced by LLM2. The functionalities of each mode are detailed below.
3.2. Live Mode
Figure 2. AuditLLM - Live Mode.
The live mode allows for a real-time interrogation of LLMs through the submission of individual queries, providing immediate insights into the model’s potential inconsistencies. The live mode comprises the following components, as illustrated in Figure 2:
  • Step 1: Users are provided with the capability to select from a predefined list of LLMs within the AuditLLM tool for auditing purposes. Currently available models include Llama 2-7B (Touvron et al., 2023), Falcon (Penedo et al., 2023), Zephyr 7B (Tunstall et al., 2023), Vicuna (Chiang et al., 2023), and Alpaca (Taori et al., 2023). More models are being added through ongoing developments.
  • Step 2: Users are required to input a query, which serves as the basis for auditing the LLM.
  • Step 3: Upon receiving user input, AuditLLM initiates the creation of customized probes aimed at evaluating the designated LLM. The prompt template for LLM1 is generated using the method outlined in (Shah, 2024), which utilizes human-in-the-loop validation and is based on relevance and diversity criteria. We are currengly using Mistral 7B LLM (Jiang et al., 2023) as our LLM1 to generate five probes that are not only closely related to the user’s initial query but also cover a wide range of inquiries, enabling a comprehensive auditing.
  • Step 4: Users are then afforded the opportunity to select from the generated probes, identifying those that most effectively facilitate the identification of inconsistencies within the audited LLM. This step is critical for narrowing down the focus of the audit to areas of potential concern or interest.
  • Step 5: In the final step, the audited LLM (LLM2 in 2) generate responses to the user-selected probes from the previous step. The semantic similarity of these responses is quantitatively determined by leveraging the BERT model (Zhang et al., 2019) via the sentence-transformer variant of MPNet (all-mpnet-base-v2) and cosine similarity, aiming to highlight sentences with semantic similarity. The culmination of this step is the presentation of a semantic similarity score, providing a metric for comparing responses and thus gauging the LLM consistency in processing and understanding queries.
Figure 2 showcases an application of this mode in the AuditLLM tool, following the steps previously outlined. AuditLLM highlights the semantic similarity between responses, applying a 60% similarity threshold. Semantically similar sentences are highlighted with their similarity scores across texts, underscoring the consistent quality of the responses to these probes. This analysis of similarities across responses allows us to identify any inconsistencies in the model’s outputs. A robust and reliable LLM should produce semantically consistent responses to similar questions, irrespective of their phrasing.
3.3. Batch Mode
The batch mode is designed for broader, more comprehensive assessments, allowing for the simultaneous submission of multiple questions for auditing the model’s inconsistencies. The components of the batch mode are:
  • Step 1: The AuditLLM tool offers users the flexibility to choose from a curated selection of the above mentioned LLMs for the purpose of auditing.
  • Step 2: Users are asked to upload a set of questions as the basis for auditing the LLM.
  • Step 3: Drawing on “Question Generation” research that emphasizes the importance of relevance and diversity in crafting effective probes (Dugan et al., 2022; Raina and Gales, 2022), users are given a choice to assign scores to these two criteria. These scores shape the creation of probes, guaranteeing relevance to user queries and a broad range of perspectives. Following (Shah, 2024), and guided by these scores, AuditLLM uses Mistral 7B LLM to generate five probes.
  • Step 4: Users are asked to provide their email addresses, facilitating the dissemination of the audit reports.
  • Step 5: AuditLLM retrieves responses from the audited LLM via tailored probes, and assesses semantic similarity using the BERT model via the sentence-transformer to identify closely related sentences. Additionally, it calculates a semantic similarity score for a quantitative evaluation of consistency. These findings, along with relevant details, are then aggregated and forwarded to the user’s email address.
Figure 3. Prompt Similarity vs Responses Similarity.
[…]
4. Conclusion
AuditLLM is a tool for auditing LLMs that is capable of uncovering inconsistencies through the exploration of varied question iterations. With its live and batch modes, AuditLLM offers a dual approach for both real-time, and extensive probing. AuditLLM, while innovative, does have its limitations. Specifically, it is designed exclusively for probing LLMs within the context of textual data and with inconsistencies in LLM’s responses as the primary mechanism to identify potential problems. Furthermore, given the substantial size of these models, loading them to generate responses can be time-intensive. We are continuing to refine this tool by adding support for more LLMs, plugging in other ways to create multiple probes (e.g., using personas), and providing more methods for assessing inconsistencies.
ACKNOWLEDGEMENTS
This work was supported by the Bill & Melinda Gates Foundation.